OpenStudy (anonymous):

multiply (square symb6 +1)(5square symb6 +1)

OpenStudy (anonymous):

\[(\sqrt{6}+1)(5\sqrt{6}+1)\]

OpenStudy (anonymous):

\[(\sqrt6+1)(5\sqrt6+1)\\ 5\sqrt{6{^2}}+6\sqrt6+1)\\ 5(6)+\sqrt{6}+1\\ 31+\sqrt{6} \]